A current of 3A flows for 50hr through a 6ohm resistor find the energy consumed by the resistor

1
Expert's answer
2022-05-20T04:53:34-0400
"W=I^2Rt\\
=3^2*6*50*3600=3.24*10^6\:\rm J"